OpenStudy (anonymous):

An ecologist wishes to mark off a circular sampling region having radius 10 m. However, the radius of the resulting region is actually a random variable R with the following pdf. f(r) = 3/4*(1-((15-r)^2)) 14≤r≤16 0 otherwise What is the expected area of the resulting circular region? (Round your answer to two decimal places.)

OpenStudy (kittiwitti1):

Is this what you're asking? \[f(r)=\frac{3}{4}(1-(15-r)^{2}),14\le r\le16\]

OpenStudy (anonymous):

Yes

OpenStudy (kittiwitti1):

Okay. So you have the radius value already, r=10, yes?

OpenStudy (anonymous):

oh yes

OpenStudy (kittiwitti1):

I believe you just need to input that value for r in the equation. Though if it's not, then I'm not sure. I haven't done Stats in awhile haha ^-^;

OpenStudy (anonymous):

I tried to put it into the equation twice but it was wrong and I've got the last chance to check answer.....

OpenStudy (kittiwitti1):

Hrm... well, you can always check answers with Mathway - http://www.mathway.com/ Cymath is also a good site to run through the steps - http://www.cymath.com/ (there's also WolframAlpha but idk if that will help much as of now)

OpenStudy (anonymous):

Oh Thanks for the help :)

OpenStudy (kittiwitti1):

Yeah, sorry > < I'm not too hot at Stats now. I need to fix that ahah @hartnn or @Directrix should be able to help maybe though, they're math pros ☺

OpenStudy (kittiwitti1):

@satellite73

OpenStudy (anonymous):

Its asking to the expected value of the pdf. i think

OpenStudy (anonymous):

\[\mathbb{E}[x] = \int_{-\infty}^\infty xf(x) dx\]

Directrix (directrix):

@kittiwitti1 Look at @Zarkon 's work here: http://openstudy.com/study#/updates/4f616c8ee4b079c5c630f224

OpenStudy (kittiwitti1):

Ah... I haven't done statistics of that advanced level. It was really just a mini introductory run-through ^-^; @Directrix My apologies @kim3584 I was not experienced enough to help you :(

Directrix (directrix):

@kim3584

OpenStudy (anonymous):

Oh I saw that already. I did not try yet so I will try it. Thanks.